OpenStudy (anonymous):

Find the vertices and foci of the hyperbola with equation. quantity x minus 3 squared divided by 81 minus the quantity of y plus 5 squared divided by 144 equals 1

OpenStudy (perl):

$$\Large \frac{(x-3)^2}{81}- \frac{ (y+5)^2}{144}=1 $$

OpenStudy (perl):

Is that the question?